# Quadratic Mode Couplings in Rotating Black Holes and Their Detectability

### v2

@@ VI Discussion
Together with Paper I, we have presented two independent frequency-domain methods for analyzing QQNMs — the complex contour technique and the hyperboloidal slicing approach. Both methods produce consistent results for $R_{++}$ of $\begin{pmatrix}220+\\ 220+\end{pmatrix}$. We have also classified the four possible coupling channels for a QQNM. Notably, the coupling between two mirrored modes (labeled by --- -) vanishes exactly. After accounting for all coupling channels in systems with reflection symmetry, our results for Kerr BHs match those from numerical simulations and time-domain fits.
We have performed a SNR survey on various QQNMs for nonprecessing binaries. Besides the known mode $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we considered two more: $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ and $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$. For an event with $M_{\rm tot}=100M_{\odot}$ and $D_{L}=410$ Mpc, the SNR for CE may reach $\mathcal{O}(8)$ for binaries with $\chi_{\rm eff}>0.6$ and $q\in(2,5)$. For SgrA*-like binaries ($M_{\rm tot}=8\times 10^{6}M_{\odot}$) at a redshift of 1, LISA’s SNR may reach $\mathcal{O}(10^{2})$. We expect LISA to detect additional QQNMs. A more detailed survey is left for future work. Future studies could also explore precessing binaries, where QQNM excitation may encode richer information about linear parent modes and progenitors’ parity. It would be interesting to explore how to decipher the information from QQNM amplitudes.

### v3

@@ VI Discussion
Together with Paper I, we have presented two independent frequency-domain methods for analyzing QQNMs — the complex contour technique and the hyperboloidal slicing approach. Both methods produce consistent results for $R_{++}$ of $\begin{pmatrix}220+\\ 220+\end{pmatrix}$. We have also classified the four possible coupling channels for a QQNM. Notably, the coupling between two mirrored modes (labeled by --- -) vanishes exactly. After accounting for all coupling channels in systems with reflection symmetry, our results for Kerr BHs match those from numerical simulations and time-domain fits, which therefore resolves the tension between Paper I and Zhu et al. [13].
The agreement between our theoretical predictions and numerical relativity results provides a compelling quantitative demonstration—one of only a few such examples (e.g. [51, 52])—of the success of BH second-order perturbation theory. Our theoretical analysis reveals that the quadratic ratios reported in [13, 14] are limited to nonprecessing binaries. By contrast, the excitation of QQNMs in precessing binaries is expected to encode richer information about parent modes and progenitors’ parity. This opens an avenue for future studies to investigate how QQNMs amplitudes can infer precession parameters.
We have also illustrated potential applications of our calculations in future ringdown data analysis. As a figure of merit, we performed an SNR survey. Our results indicate that multiple QQNMs are likely to play a nonnegligible role in BH spectroscopy with future detectors. Therefore, a robust ringdown data analysis framework will be required to account for a substantial number of linear modes and QQNMs, in order to avoid systematic biases and prevent false alarms of deviations from GR.